Fermi National Laboratory
 

Accelerator Update for February 5, 2007 - February 7, 2007

 

Monday February 5

The day shift began with Operations monitoring the Recycler stash (98.4E10), with the Antiproton Source (Pbar) stacking (128.1E10), with the Tevatron (TeV) waiting for CHL to resume normal operations, with feeders 36 and 37 off, with Booster dealing with a vacuum problem, with the Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beams.

A TeV expert made some repairs to the TeV sector C17 separator at 8:13 AM.

The High Voltage Coordinator reported at 8:33 AM that feeder 36 had tripped off due to an oil leak in a PS4 transformer. The HV Coordinator turned off feeder 37.

Cryo system technicians at CHL reported at 9:04 AM that CHL was back up and running.

Water technicians reported at 9:26 AM that the Meson LCW system was up and running.

At 9:51 AM, Operations tried, unsuccessfully, to tune beam to Switchyard (SY). Operators discovered that the MI-52 septa high voltage was off. They turned it on and began sending beam to Pbar, NuMI, and Meson.

A Recycler kicker expert reported at 9:52 AM that a kicker (R:KPS2A) had gotten too cold. He believed that warming the magnet would clear the alarm.

At 9:58 AM, Operations reported that a Pbar Accumulator kicker (A:IKIK) wasn’t working due to an arcing cable. An expert began repairs.

Operations put the TeV through a dry squeeze at 11:12 AM.

The HV Coordinator reported at 11:23 AM that he had energized feeder 36.

The kicker expert reported that A:IKIK was repaired. Pbar resumed stacking at 11:42 AM.

Operations put the TeV into shot setup at 12:18 PM.

At 1:17 PM, a Linac Klystron RF station (KRF1) tripped off on a LCW flow fault. Operators had the station reset within a couple of minutes.

Operations began preparing MI at 2:10 PM for an access to investigate the Recycler kicker cooling level.

At 2:20 PM, KRF3 tripped off due to a rack over temperature alarm. Experts investigated and found the rack-cooling fan had failed.

At 2:38 PM, in the process of preparing MI for access, an operator inadvertently pressed the TeV “test/reset” button for the power supply loop controller. There was no quench, but the store was aborted.

Operations prepared D0 for an access at 3:33 PM.

Recycler experts reported at 3:41 PM that the kicker (KPS2A) was fine. The cooling fluid was okay. Operations successfully reset the magnet.

At 3:52 PM, Operations reported that a Booster kicker (B:MKS07) had tripped off due to a fan fault. An expert investigated and had it back on within twenty minutes.

D0 came out of access at 4:23 PM.

At 4:52 PM, a MI quadrupole power supply (QT101) was not regulating properly. Operators investigated and turned off the supply.

At 5:21 PM, MI experts reported that they could not tune around QT101. Operators reported at 5:54 PM that they had replaced the controller. MI resumed sending beam.

TeV experts began a study period around 6 PM.

At 7:19 PM, Operations reported that BRF10 had tripped off a couple of times. This seemed to be due to a problem with BRF12.

At 8:34 PM, BRF6 tripped off. It reset.

Operations put the TeV into shot setup at 9:26 PM.

At 10:08 PM, a Booster expert reported that BRF12 had no gap envelope. An RF expert reported that there was a problem with its modulator. Another expert was called in.

At 10:38 PM, Operations reported that due to the loss of Booster RF stations, Booster intensity was down.

At 10:59 PM, BRF6 tripped off. It reset.

Operations established store 5215 at 11:53 PM with an initial luminosity of 86.8E30.

Operations reported at midnight that the Recycler had made one stash during the evening shift.

Tuesday February 6

The midnight shift began with Operations monitoring store 5215 (77.13E30) and the stash (65.7E10), with Pbar stacking stack (8.2E10), and with NuMI taking beam.

At 2:44 AM, BRF6 tripped off. It reset.

Operations reported at 8:15 AM that the Recycler had stashed three times during the midnight shift.

At 8:51 AM, Operations reported Controls problems; they couldn’t load pages or ACNET applications. Operations contacted experts.

An RF expert reported at 9:22 AM that he had BRF12 back up and running.

Operations reported at 9:36 AM that they had established beam to SY.

Operations reported at 9:56 AM that the Controls problem went away. Experts were still investigating. Much later in the day, a Controls expert reported that a new release of the database API used by Java was repeatedly trying to connect to the database. Experts backed out of the API and restarted the Java service.

At 10:26 AM, BRF6 tripped off for the first of six times in an hour and a half. It reset each time.

At 10:42 AM, BRF12 tripped off for the first of two times in ten minutes and wouldn’t reset. Operators bypassed the station. Experts investigated.

At 11:34 AM, Operations reported a LCW supply problem at the F1 service building. Water experts found the pond pump running, but with no flow. They started up the backup pump.

At 12:04 PM, a Linac Klystron RF station (KRF5) tripped off for the first of five times in ten minutes. It eventually reset after each trip.

At 12:40 PM, BRF6 tripped off. It reset okay.

Operations made the Meson enclosures M03 & 4 ready for access at 1:20 PM.

At 1:31 PM, BRF6 tripped and wouldn’t reset. Operators bypassed the station and contacted experts.

Operations terminated store 5215 at 2:09 PM.

Operations put the TeV into shot setup at 2:26 PM.

The Meson access to M03 & 4 ended at 2:51 PM.

At 2:55 PM, KRF5 tripped off. A Linac expert had the station back up within six minutes.

At 3:09 PM, Operations reported that the MI SBD wasn’t working. Operations contacted experts.

A Pbar expert reported at 3:35 PM that they were testing a new lithium lens and transformer in the AP0 test stand.

Operations reported at 3:58 PM that support techs were replacing a bad Recycler trim regulator (R:H616). Also, Pelletron experts were accessing their pit to investigate an electron beam energy drift. They replaced some bad cables.

Operations reported at 4:22 PM that the Recycler had stashed two times during the day shift.

Operations established store 5216 at 4:25 PM with an initial luminosity of 208.8E30.

At 4:58 PM, the NuMI beam permit tripped off due to a bad beam position. It reset okay.

SY experts finished using the Meson beam at 6:36 PM.

At 7:55 PM, BRF3 tripped off and wouldn’t reset. Operators bypassed the station.

At 10:29 PM, BRF13 & 14 tripped off due to a vacuum pump trip. Everything reset okay.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday February 7

The midnight shift began with Operations monitoring store 5216 (78.1E30) and the stash (117.4E10), with Pbar stacking (44.3E10), with the MI, Recycler, and Linac operating normally, with the Booster intensity low due to RF problems, with NuMI taking beam, with MiniBooNE off, and with CDF and D0 using the colliding beams.

Operations turned off the Booster at 6 AM in preparation for a tunnel access to repair RF.

Operations turned off the Pelletron at 6:07 AM because it had tripped off too many times.

Operations put Pbar in standby at 6:11 AM.

RF experts accessed the Booster at 7:03 AM.

Operations prepared the Linac for an access at 7:20 AM.

Operations reported at 8:17 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plan for today is to complete the Booster RF repairs by noon. Pbar and Linac experts will access their tunnels for maintenance work during this time.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/07/2007    email Fermilab